OpenStudy (anonymous):

When a 7.00-g sample of KCl is dissolved in water in a calorimeter that has a total heat capacity of 5.383 kJ·K–1, the temperature decreases by 0.300 K. Calculate the molar heat of solution of KCl.

OpenStudy (photon336):

I think that this would have to be an endothermic reaction, because energy is absorbed. given by the fact hat the temperature decreases (tf < ti) meaning that heat was absorbed from the surroundings. \[7.0g KCl*(\frac{ mol }{ 75.0g }) = 0.09 mol, KCl\] \[q = mC \delta t \] not sure: \[\frac{ 5,383J }{ k (0.09) mol } = 57.9 kj*k ^{-1}mol ^{-1}\] q = 5.383jk*K^-1

OpenStudy (photon336):

@Cuanchi

OpenStudy (photon336):

@Rushwr C = specific heat, while q = heat capacity right? so all you need to do is put q/moles that's what I thought for this question.

OpenStudy (anonymous):

57.9 is wrong.

OpenStudy (anonymous):

122 is also wrong.

OpenStudy (photon336):

@Cuanchi @Zale101

OpenStudy (photon336):

had a feeling something was wrong, whoever gets the right answer to this please tag me. I want to see how this is done. @aaronq

OpenStudy (anonymous):

Nope, I've tried so many methods and all of them have been wrong so far haha. It should be in kj/mol.

OpenStudy (anonymous):

omg. it was as simple at (-.3)(5.383) = 1.61 then divide by (7g/74.5mol) = 17.2

OpenStudy (anonymous):

17.2 kj/mol is the answer!

OpenStudy (anonymous):

Heat absorbed by calorimeter = (heat capacity of calorimeter)(Tf - Ti) then divide by moles to get into the correct units

OpenStudy (anonymous):

thanks for all the help!